Two coin's are tossed. Let H be the number of Tails that occur. Determine the values of the random variable H.

1
Expert's answer
2022-02-22T15:04:06-0500

This is a binomial distribution with "n=2,p=\frac{1}{2}."

"P(H=0)=C^0_2(\frac{1}{2})^0(\frac{1}{2})^2=\frac{1}{4}."

"P(H=1)=C^1_2(\frac{1}{2})^1(\frac{1}{2})^1=\frac{1}{2}."

"P(H=2)=C^2_2(\frac{1}{2})^2(\frac{1}{2})^0=\frac{1}{4}."